Certain non-metal X forms two oxides I and II . The mass percentage of oxygen in I(X4 O6) is 43.7% which is the same  as that of X in the 2nd oxide.Find the formula of 2nd oxide. (Based on law of multiple proportions...... pls xplain each step carefully) 

Grade:11

1 Answers

Chetan Mandayam Nayakar
12 years ago

Dear Ajay,

let 'm' be atomic mass of X. let formula of second oxide be XuOv

mass %age of oxygen in I=96/(96+4m)=24/(m+24)=0.437,m+24=24/0.437,m=31(approx.)

31u/(31u +16v)=0.437,31u/0.437=31u+16v

31/0.437=31+16(v/u)

v/u=2.5,therefore II is X2O5

Please feel free to post as many doubts on our discussion forum as you can. If you find any question
Difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We
are all IITians and here to help you in your IIT JEE preparation.


All the best Ajay!!!

 


Regards,

Askiitians Experts

CHETAN MANDAYAM NAYAKAR